12. Investors require a 15% rate of return on Levine Companys stock (that is, r, = 15%). a. What is the value if the previouI have part A but need help with B and C
Answer #1

12 b. D0=2
1) Price at growth Rate of 15%=D0*(1+g)/(R-g) =2*(1+15%)/(15%-15%) =Cannot be determined
2) 1) Price at growth Rate of 20%=D0*(1+g)/(R-g) =2*(1+20%)/(15%-20%) =-48

These are not reasonable results as when growth is more than required rate price cannot be determined or price is negative.
This is because when required rate is less than or equal to growth rate then the results doesnot make sense.

c. Growth can be greater than required rate for short period
but it is not reasonable for growth rate to be greater than required rate for indefinite period.